# Thermal state preparation of the SYK model using a variational quantum algorithm

### v1

@@ I Introduction
Due to its role in holography, a detailed study of both pure and thermal (or Gibbs) states of the SYK model is necessary. The ground (pure) state of the SYK model exhibits a volume law entanglement, which makes them classically hard to prepare [6]. In addition, a defining property of the SYK model is that thermal four-point correlators, in particular out-of-time order correlators (OTOCs), saturate the exponent that measures the exponential growth of such correlators in the large $N$ and low-temperature limit [7]. To measure these OTOCs, we have to prepare the relevant thermal state, which will be the focus of this work. While preparing low-energy states, which have been studied using large-scale classical simulations, our work aims to perform benchmark calculations using hybrid quantum/classical algorithms. The recent progress in quantum technologies and algorithms has led to various explorations of quantum computing in the context of fundamental physics [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. In elementary particle and nuclear physics, quantum computing promises to enable calculations that lie beyond the reach of classical simulations. In this work, we carry out hybrid classical/quantum simulations using variational quantum algorithms and small-scale benchmark simulations on quantum hardware. In recent years, several studies of the SYK model have been carried out using quantum computing methods [27] such as teleportation [28], computation of bosonic correlation functions [29], real-time dynamics [30, 31], and the ground state preparation of the (coupled) SYK model [32, 33].
@@ II The dense and sparse SYK model
One of the challenges of the full or dense SYK model is that the number of terms in the Hamiltonian grows as $\sim N^{4}$. To reduce the computational cost of classical or quantum simulations of the SYK model, a sparsified version was introduced in Ref. [62, 63]. See also Refs. [32, 62, 31]. In the large-$N$ limit, the SYK model has $\sim N^{4}/24$ terms and an average connectivity per Majorana fermion of $k\sim N^{3}/24$. Any SYK-type model with $k\ll N^{3}/24$ is referred to as a sparse SYK model. The value of $k$ can be chosen as small as possible to reduce the computational cost but sufficiently large to retain a holographic interpretation, see Ref. [65]. Following Ref. [63], the sparsified SYK model is obtained by modifying the Hamiltonian in Eq. (1) as
@@
The variance of the couplings $J_{ijkl}$ is more significant compared to Eq. (2), which ensures that the energy scales are comparable to the original SYK model. The probability with which the terms are retained is related to the average connectivity between fermions $k$ as $p\approx 24k/N^{3}$. Therefore, the variance of the sparse SYK model is $\approx 3!\mathcal{J}^{2}/24k$ and independent of the number of fermions $N$ for a given value of $k$. The dense or full SYK model corresponds to $p=1$. It was shown in Ref. [64] that until one goes below the percolation limit, the Lyapunov exponent does not depend strongly on the sparsity of the model. This work considers the sparse SYK model with a fixed sparsity of $k=8.7$.
@@ III Variational thermal quantum algorithm
Different than in Eq. (8), $p_{i}({\bm{\theta}})=|a_{i}({\bm{\theta}})|^{2}$ corresponds to the probability of the $i^{\rm th}$ basis state in the ensemble. To calculate the free energy in Eq. (7), we approximate the von Neumann ${\cal S}$ with the Shannon entropy $S$. Note that the von Neumann entropy would require access to the entire density matrix, whereas the Shannon entropy can be calculated from its diagonal entries. This simplification can sometimes result in a less accurate construction of the thermal state. The measurement outcomes in the computational basis of VQC1 are now taken as the initial state for the second part of the algorithm. We denote the unitary operation of the set $\left\{\text{VQC}_{2}\right\}$ by $U_{2}({\bm{\phi}})$, where ${\bm{\phi}}$ denotes the variational parameters. Different than for VQC1, we start with a Hamiltonian-based ansatz [74], i.e. the variational circuit consists of gates that appear in the trotterized time evolution operator associated with the Hamiltonian given in Eq. (1).
@@ A.1 Choice of the cost function
In the main text, we used the free energy as the cost function to construct an approximation to the thermal state of the system at some inverse temperature $\beta$. This can be seen by considering the relative entropy between two density matrices, $\rho_{p}$ and $\rho_{\beta}$, where the former denotes the approximation we prepared and the latter the true thermal density matrix. The variationally prepared thermal state $\rho_{p}$ is close to the actual thermal state if the relative entropy $S(\rho_{p}||\rho_{\beta}$) given in (16), which is non-negative due to Gibbs inequality is close to zero. The relative entropy is minimized when $F_{p}$, i.e., the prepared free energy, is equal to the actual free energy. The relative entropy is given by $\beta(F_{p}-F_{\beta})$ i.e., (17).
@@
For completeness, we derive these results in the following. First, we show that relative entropy is lower bound by zero. Then, we show that the trial free energy is always larger than the actual free energy. To show the first result, consider two density matrices given by $\rho=\sum_{n}p_{n}\ket{n}\bra{n}$ | and $\rho^{\prime}=\sum_{n}p_{n}^{\prime}\ket{n}\bra{n}$ | respectively. We find
$\rho\log(\rho^{\prime})-\rho\log(\rho)=\sum_{n}\big{[}p_{n}\log(p_{n}^{\prime})-p_{n}\log(p_{n})\big{]}\ket{n}\bra{n}\,.$ | . (14)
Taking the trace on both sides and using $\log(x)\leq x-1$, we obtain
$\rm{Tr}[\rho\log(\rho)]-\rm{Tr}[\rho\log(\rho^{\prime})]$ $=\sum_{n}(p_{n}\log(p_{n})-p_{n}\log(p_{n}^{\prime}))$
$=\sum_{n}p_{n}\log\Big{(}\frac{p_{n}}{p_{n}^{\prime}}\Big{)}$
@@
After rewriting the relative entropy in terms of thermodynamic observables, we find
$S(\rho_{p}||\rho_{\beta})$ $={\rm Tr}[\rho_{p}\log(\rho_{p})]-{\rm Tr}[\rho_{p}\log(\rho_{\beta})]$ (16)
@@
Note that $F_{p}\geq F_{\beta}$ is sometimes referred to as the Gibbs-Bogoliubov inequality. This inequality states that at any finite temperature $T$, the free energy of the system is always smaller than one calculated using some trial state. This inequality is the finite-temperature version of the Rayleigh-Ritz result. To construct the thermal state, one can also maximize the von Neumann entropy. However, it is computationally easier to minimize the free energy, as done in this work.

### v2

@@ I Introduction
Due to its role in holography, a detailed study of both pure and thermal (or Gibbs) states of the SYK model is necessary. The ground (pure) state of the SYK model exhibits a volume law entanglement of the entanglement entropy with a known coefficient [6], which makes them classically hard to prepare [7]. In addition, a defining property of the SYK model is that thermal four-point correlators, in particular out-of-time order correlators (OTOCs), saturate the exponent that measures the exponential growth of such correlators in the large $N$ and low-temperature limit [8]. To measure these OTOCs, we have to prepare the relevant thermal state, which will be the focus of this work. While preparing low-energy states, which have been studied using large-scale classical simulations, our work aims to perform benchmark calculations using hybrid quantum/classical algorithms. The recent progress in quantum technologies and algorithms has led to various explorations of quantum computing in the context of fundamental physics [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In elementary particle and nuclear physics, quantum computing promises to enable calculations that lie beyond the reach of classical simulations. In this work, we carry out hybrid classical/quantum simulations using variational quantum algorithms and small-scale benchmark simulations on quantum hardware. In recent years, several studies of the SYK model have been carried out using quantum computing methods [28] such as teleportation [29], computation of bosonic correlation functions [30], real-time dynamics [31, 32], and the ground state preparation of the (coupled) SYK model [33, 34].
@@ II The dense and sparse SYK model
One of the challenges of the full or dense SYK model is that the number of terms in the Hamiltonian grows as $\sim N^{4}$. To reduce the computational cost of classical or quantum simulations of the SYK model, a sparsified version was introduced in Ref. [63, 64]. See also Refs. [33, 63, 32]. In the large-$N$ limit, the SYK model has (${{N}\choose{4}}\sim N^{4}/24$ terms. The number of terms in the Hamiltonian per Majorana fermion is $k=N^{3}/24$, which is also called the average connectivity of the corresponding hypergraph. When $k\ll N^{3}/24$, the resulting Hamiltonian is referred to as the sparse SYK model. The value of $k$ can be chosen as small as possible to reduce the computational cost but sufficiently large to retain a holographic interpretation, see Ref. [66]. In this work, we consider the sparse SYK model with a fixed sparsity of $k=8.7$. Following Ref. [64], the sparsified SYK model is obtained by modifying the Hamiltonian in Eq. (1) as
@@
The variance of the couplings $J_{ijkl}$ is more significant compared to Eq. (2), which ensures that the energy scales are comparable to the original SYK model. The probability with which the terms are retained is related to the average connectivity between fermions $k$ as $p\approx 24k/N^{3}$. Therefore, the variance of the sparse SYK model is $\approx 3!\mathcal{J}^{2}/24k$ and independent of the number of fermions $N$ for a given value of $k$. The dense or full SYK model corresponds to $p=1$. It was shown in Ref. [65] that until one goes below the percolation limit, the Lyapunov exponent does not depend strongly on the sparsity of the model.
@@ III Variational thermal quantum algorithm
Different than in Eq. (8), $p_{i}({\bm{\theta}})=|a_{i}({\bm{\theta}})|^{2}$ corresponds to the probability of the $i^{\rm th}$ basis state in the ensemble. To calculate the free energy in Eq. (7), we approximate the von Neumann ${\cal S}$ with the Shannon entropy $S$. Note that the von Neumann entropy would require access to the entire density matrix, whereas the Shannon entropy can be calculated from its diagonal entries. The Shannon entropy is equal to von Neumann entropy only in the Schmidt basis. This simplification can sometimes result in a less accurate construction of the thermal state. However, even for quantum computers an explicit computation of von Neumann entropy is possibly not efficient. An optimal thermal state algorithm potentially has to converge to a convex mixture, which is the closest approximation to the Schmidt basis where the Shannon and von Neumann entropy agree. The measurement outcomes in the computational basis of VQC1 are now taken as the initial state for the second part of the algorithm. We denote the unitary operation of the set $\left\{\text{VQC}_{2}\right\}$ by $U_{2}({\bm{\phi}})$, where ${\bm{\phi}}$ denotes the variational parameters. Different than for VQC1, we start with a Hamiltonian-based ansatz [75], i.e. the variational circuit consists of gates that appear in the Trotterized time evolution operator associated with the Hamiltonian given in Eq. (1).
@@ A.1 Choice of the cost function
In the main text, we used the free energy as the cost function to construct an approximation to the thermal state of the system at some inverse temperature $\beta$. This can be seen by considering the relative entropy between two density matrices, $\rho_{p}$ and $\rho_{\beta}$, where the former denotes the approximation we prepared and the latter the true thermal density matrix. The variationally prepared thermal state $\rho_{p}$ is close to the actual thermal state if the relative entropy $S(\rho_{p}|\rho_{\beta}$) is close to zero. The relative entropy is minimized when $F_{p}$, i.e., the prepared free energy, is equal to the actual free energy.
@@
For completeness, we derive these results. First, we show that relative entropy is lower bounded by zero. Then, we relate it to the prepared and exact free energy. To show the first result, consider two density matrices given by $\rho=\sum_{i}p_{i}\ket{i}\bra{i}$ | and $\rho^{\prime}=\sum_{j}q_{j}\ket{j}\bra{j}$ | respectively. We find
$S(\rho|\rho^{\prime})$ $=\rm{Tr}(\rho\log\rho)-\rm{Tr}(\rho\log\rho^{\prime})$
$=\sum_{i}p_{i}\log p_{i}-\sum_{i}\langle i|\rho\log\rho^{\prime}|i\rangle$
$=\sum_{i}p_{i}\log p_{i}-\sum_{i}\sum_{j}p_{i}\langle i|j\rangle\langle j|i\rangle\log q_{j}$
$=\sum_{i}p_{i}\Big{(}\log p_{i}-\sum_{j}\langle i|j\rangle\langle j|i\rangle\log q_{j}\Big{)}$ (14)
Since the logarithm is a concave function, we have $\log(a_{i})\geq\sum_{j}\langle i|j\rangle\langle j|i\rangle\log(q_{j})$ where $a_{i}=\sum_{j}\langle i|j\rangle\langle j|i\rangle q_{j}$. Using this in (14), we find
$S(\rho|\rho^{\prime})$ $\geq\sum_{i}p_{i}\Big{(}\log p_{i}-\log a_{i}\Big{)}$
$\geq-\sum_{i}p_{i}\log\Big{(}\frac{a_{i}}{p_{i}}\Big{)}\,.$ (15)
Using $1-x\leq-\log(x)$ for $x>0$, we obtain
$S(\rho|\rho^{\prime})$ $\geq-\sum_{i}p_{i}\log\Big{(}\frac{a_{i}}{p_{i}}\Big{)}$
$\geq-\sum_{i}p_{i}\Big{(}1-\frac{a_{i}}{p_{i}}\Big{)}$
@@
With this result that the relative entropy (also known as ‘relative entropy of entanglement’) is non-negative, we can express it in terms of thermodynamical observables and find:
$S(\rho_{p}|\rho_{\beta})$ $={\rm Tr}[\rho_{p}\log(\rho_{p})]-{\rm Tr}[\rho_{p}\log(\rho_{\beta})]$
@@
Note that $F_{p}\geq F_{\beta}$ is sometimes referred to as the Gibbs-Bogoliubov inequality. This inequality states that at any finite temperature $T$, the free energy of the system is always smaller than the one calculated using some trial state. This inequality is the finite-temperature version of the Rayleigh-Ritz result. To construct the thermal state, one can also maximize the von Neumann entropy. However, it is computationally easier to minimize the free energy, as done in this work.